Determine the oxidation state for each of the elements below.

The oxidation state of

...

phosphorus

...

in

...

phosphorous tribromide
PBr3

...

is

...

The oxidation state of

carbon

in

carbon monoxide
CO

is

The oxidation state of

oxygen

in

potassium peroxide
K2O2

is
Answer #1

1)

Oxidation number of Br = -1

lets the oxidation number of P be x

use:

3* oxidation number (Br) + 1* oxidation number (P) = net charge

3*(-1)+1* x = 0

-3 + 1 * x = 0

x = 3

So oxidation number of P = +3

Answer: +3

2)

Oxidation number of O = -2

lets the oxidation number of C be x

use:

1* oxidation number (O) + 1* oxidation number (C) = net charge

1*(-2)+1* x = 0

-2 + 1 * x = 0

x = 2

So oxidation number of C = +2

Answer: +2

3)

This is ionic compound of K+ and O2-2

lets the oxidation number of O be x

use:

1* oxidation number (O) = net charge

2* x = -2

2 * x = -2

x = -1

So oxidation number of O = -1

Answer: -1